In triangle​ ABC, the angles A and B have the same​ measure, while the measure of angle C is 63 degrees larger than the measure of each of A and B. What are the measures of the three​ angles?
Thank you for your help!

Answer by math_helper(2461)


The three angles must add to 180 degrees.
We can write angle B as "A" because they are equal, giving us:
2A + (A+63) = 180
3A + 63 = 180
3A = 117
A = 39 —> B = 39, C = 39+63 = 102

Ans: The angles A & B are each 39 degrees, while angle C measures 102 degrees.

Check: 39+39+102 = 78+102 = 180 (ok)